OpenStudy (anonymous):

13 + w/7 = –18 (thirteen plus w over seven equals negative eighteen.)

OpenStudy (compassionate):

13 + w/7 = -18 -13 -13 w/7 = -31 w/7 * 7 = -31 * 7 w = -31 * 7 Do the last part

OpenStudy (anonymous):

-217?

OpenStudy (compassionate):

Yup.